OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). x + y = 15 4x + 3y = 38

OpenStudy (anonymous):

3*(1) - (2) 3x+3y=45 => 3x-4x +3y-3y=45-38 -x=7 x=-7 -7+y=15 y=22

OpenStudy (anonymous):

x=15-y 4(15-y)+3y=38 6-4y+3y=38 -y=38-60 y=22 22+x=15 x+15-22 x=-7 (-7,22)

OpenStudy (anonymous):

Take the bottom from 4* the top -> y=22 and sub back for x.

OpenStudy (anonymous):

y = 15 - x plug this into second equn. 4x + 45- 3x = 38 x = -7 y = 15 - - 7 = 22 answer x,y = -7,22

OpenStudy (anonymous):

-7,22 thats wrong

OpenStudy (anonymous):

(-7,22) better in brackets?

OpenStudy (anonymous):

-7,22 thats wrong -7 + 22 = 15 => correct! 4*(-7) + 3*22 = -28 + 66 = 38 => correct!

OpenStudy (llort):

(-7, 22)

OpenStudy (anonymous):

x+y=15 4x+3y=38 4x+4y=60 4x+3y=38 then y=22 x+22=15 x=15-22=-7 ans,x y=-7,22